Nitrogen trifluoride gas reacts with steam to produce the gases HF, NO, and NO2.

a. what volume of nitrogen oxide is formed when 5.22L of nitrogen trifluoride are made to react with 5.22L of steam? Assume 100% yields and constant temperature and pressure condition throughout the reaction.



1
Expert's answer
2021-01-12T06:59:07-0500

"2 NF_{3(g)}+ 3 H_2O_{(g)} \to 6 HF_{(g)} + NO_{(g)} + NO_{2(g)}"


5.22L of NF3 = "\dfrac{5.22}{22.4}" moles = 0.233 moles


5.22L of H2O = "\dfrac{5.22}{22.4}" moles = 0.233 moles


from the chemical reaction above, 2 moles of NF3 react with 3 moles of Steam.

Since, in this reaction 0.233 moles of NF3 react with 0.233 moles of Steam,

the limiting reagent is Steam, and the excess reagent is NF3.


from the reaction above, 3 moles of water produces 1 mole of NF3,

5.22L of water will therefore produce 1.74L ("\dfrac{5.22×1}{3}" L) of NF3.